OpenStudy (anonymous):

The age of a father is 2 less than 7 times the age of his son. In 3 years, the sum of their ages will be 52. If the son’s present age is s years, which equation models this situation?

OpenStudy (anonymous):

(s+3)+[(7s−2)+3]=52 3s+3+3(7s−2)=52 s+(7s−2)=52 (s−3)+(7s−2)−3=52